Andy wanted to go shopping, but didn't have very much money for what he wanted. He did, however, spend $17 on a shirt, 1/4 of the remaining amount on shoes, $741 on a jacket, $2000 on a horse, 1/4 of the remaining amount on a saddle, 4/5 of rest of his money on horse grain, and $150 for one month's board for the horse. He was left with nothing. How much money did he spend?
Answer by jorel555(1290)

Let n be the amount Andy had before he bought the horse grain. His last expense was 150. 150 represents 1/5 of what he had left. So, working backward:
n/5=150
n=750.
Before that, he spent 1/4 of what he had left on a saddle, so 750 represents 3/4 of what he had, so
3n/4=750
n=750*4/3=1000
Add 2741 for the horse and the jacket, and we're up to 3741--
3n/4=3741
n=4988
Add 17 for the shirt, and we have $5005 as the total amount of money Andy spent!!!!!!